OpenStudy (ineedhelpfast12):

solve the linear system using substitution. 4x-7y=10 y=x-7

OpenStudy (abb0t):

http://www.wolframalpha.com/

OpenStudy (abb0t):

They do step by step instructions on how to solve this! Far better instructions than whoever is going to help you.

OpenStudy (ineedhelpfast12):

ok

pooja195 (pooja195):

4x-7y=10 <---e1 y=x-7 <---- e2 y value is given so plug it into e1 \[\huge~\rm~4x−7(x−7)=10 \] solve for x

OpenStudy (abb0t):

you're welcome.

OpenStudy (ineedhelpfast12):

x=-13? y=-20?

pooja195 (pooja195):

-20 is wrong

OpenStudy (ineedhelpfast12):

but i got -13 for my x value

pooja195 (pooja195):

didn't see the negative sign....try again

pooja195 (pooja195):

4x-7y=10 <---e1 y=x-7 <---- e2 y value is given so plug it into e1 \[\huge~\rm~4x−7(x−7)=10\] solve for x

OpenStudy (ineedhelpfast12):

but -7*-7=49

pooja195 (pooja195):

4x-7x+49=10 go on from here

OpenStudy (ineedhelpfast12):

-3x=-39

OpenStudy (ineedhelpfast12):

oh i see. -x=-13, but I want the -x to be positive, so x=13

pooja195 (pooja195):

Correct :-) now find the y value

OpenStudy (ineedhelpfast12):

y=6

pooja195 (pooja195):

Perfect! :)

TheSmartOne (thesmartone):

Thank you @abb0t

OpenStudy (abb0t):

You're very welcome @TheSmartOne

TheSmartOne (thesmartone):

I got a 100% on my test, I can never repay you back for your service.

OpenStudy (abb0t):